# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v2

@@ 5.2 Adjudicator
Social welfare functions select an alternative based on input preferences of voters. The pros and cons of individual social welfare functions have been extensively researched and debated in social choice (Arrow, Sen, and Suzumura 2010; Rawls 2017). We make use of cardinal social welfare functions (Keeney and Kirkwood 1975) which take as input our alignment scores $(s_{i}(R))_{i\in[\ell],R\in\mathcal{R}}$ and output the winning reward function. We consider the three arguably most popular social welfare functions:
Utilitarian
@@
For each preference clause $p_{i}\in P$, we compute a reward function $R_{i}$ aligned with $p_{i}$ by casting it as a singular prompt to the DLM pipeline (see Section C.2). For each $R\in\mathcal{R}$, we compute as $s_{i}(R)$ the expected reward according to reward function $R_{i}$ produced by policy $\Pi(R)$ (again, we approximate this quantity by running multiple simulations). $s_{i}(R)$ quantifies the quality of the policy induced by the candidate reward function $R$ from the perspective of $p_{i}$ (as captured by $R_{i}$). As the scale of the reward functions can vary significantly among preference clauses, we normalize the scores by the performance of the default policy, i.e., we compute $\frac{s_{i}(R)-s_{i}(R^{*})}{s_{i}(R^{*})}$.
@@
The Simulator Scorer Model assumes access to reward functions capturing individual preference clauses well. If no well-aligned reward functions can be obtained, the performance of SCLM-SIM can deteriorate because it can become very noisy. Another disadvantage of SCLM-SIM is that the scores in SCLM-SIM are all computed via simulation, which can become computationally quite costly. Motivated by this, we propose a quicker and more flexible LLM-based approach, where we prompt an LLM to rate the alignment of a candidate reward function with a preference clause. In particular, for each $R\in\mathcal{R}$ and $p_{i}\in P$, we use a prompt that includes $R$, $p_{i}$, and the utility feature distributions produced by policy $\Pi(R)$. We ask the LLM to rank how well $R$ aligns with the preference clause $p_{i}$ on a scale from 1 to 5 (see Appendix E for prompt texts).
@@
Aligning reward functions to a prioritization prompt might cause (unintended) utility shifts in other features (e.g., due to feature correlations, shifting utility to low-income beneficiaries might shift it away from more educated ones). See Appendix B for a concrete example. SCLM offers users the option to explicitly prevent these shifts by adding additional clauses (“objectives”) to the prompt: Given a prompt $P$ (e.g., the prompt from Example 4.1), for each feature not referenced in the prompt, the user can add a new preference clause requesting a minimum shift in the utility distribution of this feature (e.g., for Example 4.1 they could add “do not change the utility distribution for feature $C$”). To compute the alignment score $s_{i}(R)$ between a reward function $R$ and a clause $p_{i}$=“minimize utility shift for feature $X$”, we compare feature $X$’s utility distribution under the default policy with its utility distribution under the policy $\Pi(R)$. Specifically, we quantify the difference using the Earth mover’s distance (EMD) between the two distributions. Afterward, we apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$, which are input to the social welfare function (along with the alignment scores for the other clauses).
@@
6 Experiments
We describe our testset (Section 6.1), the compared methods (Section 6.2), and our experimental results both for dealing with composite prioritization prompts (Section 6.3) and additionally minimizing unintended side effects (Section 6.4). Following the work of Behari et al. (2024), which constitutes our most important baseline, we use Gemini Pro (Anil et al. 2023) as the LLM in our experiments.
6.1 Dataset Description
Figure 3: Results comparing the quality of reward design methods for composite prioritization prompts. Results are averaged across $180=12\cdot 15$ values: $12$ composite prompts on $15$ RMAB instances (from $3$ datasets). Error bars represent std-error.
@@ 6.1 Dataset Description
We create three synthetic datasets, each with three features ($A$, $B$, and $C$). For some agent and feature, we randomly sample the agent’s feature value between 00 and $1$. Arm’s passive transition probabilities ($P(s,a=0,s^{\prime})$ for $s,s^{\prime}\in[0,1]$) are sampled uniformly between 00 and $1$. Our three datasets differ in how active transition probabilities are sampled. For each dataset, we set a three-dimensional weight vector $W\in[0,1]^{3}$ specifying how much each feature impacts the effect of applying an active action (see Section C.3 for details). For each agent, let $f$ denote their feature values; we sample their active transition probabilities for $s,s^{\prime}\in[0,1]$ as $P(s,a=1,s^{\prime})=P(s,a=0,s^{\prime})+\delta,\;\delta\sim\mathcal{N}(\Delta,\sigma),$ where $\Delta=W\cdot f^{T}$ and $\sigma$ is a hyperparameter. Subsequently, we discretize the values of all features into $5$ different equal-sized buckets to be consistent with the real-world domain. For each dataset, we sample $5$ instances according to the above procedure with $N=2100$ arms, a budget of $B=210$, and a time horizon of $T=12$ (to replicate the setup of the real-world domain described below where roughly $3\%-10\%$ of beneficiaries receive a resource every week for up to three months).
@@
ARMMAN (2024) is a non-profit in India that operates large-scale Maternal and Child Care Mobile Health programs for underserved communities. One of their programs disseminates critical health information via weekly automated voice messages. The goal of the NGO is to maximize beneficiaries’ engagement, i.e., the number of messages they listen to. A limited number of beneficiaries are called by health workers every week to boost engagement. The problem of planning which beneficiaries to call has been modeled and solved as an RMAB, where the good/bad state corresponds to a high/low weekly engagement of the beneficiary. We use anonymized data from a quality improvement study conducted in January 2022 (Verma et al. 2023). For each beneficiary, we have access to their income, education, and age level, which we use as our three features. Beneficiaries’ historic listenership values are used to estimate their transition probabilities under the passive action (Mate et al. 2022). One problem in estimating transition probabilities under the active action is that due to the limited number of service calls made, such transitions are rare. Thus, active transition probability estimates are noisy. To alleviate this issue, we use the features and passive transition probabilities from ARMMAN together with the procedure to sample active transitions described in the synthetic domain. Again, we create datasets for three different weight vectors $W$. Each dataset consists of five sampled RMAB instances with $N=2100$ arms, a budget of $B=210$ and a time horizon of $T=12$.
@@
6.2 Models & Baselines
We analyze four different variants of SCLM differing in the used social welfare function (Utilitarian or Egalitarian) and scorer model (Simulator or LLM), e.g., we denote as SCLM-SIM-Egalitarian SCLM with the Simulator Scorer Model and the Egalitarian social welfare function. [^1] We also ran our experiments for the Nash welfare function. Its performance was very similar, yet slightly inferior to the Utilitarian welfare function in all relevant evaluation dimensions. In our generator, we generate $4$ candidate reward functions in each step and run $5$ iterations to generate a total of $20$ candidate reward functions. In addition, we consider several LLM-focused baselines (see Appendix E for detailed descriptions):
LLM-Zeroshot
@@ 6.2 Models & Baselines
DLM
This baseline implements the Decision-Language Model by Behari et al. (2024) (see Section C.2).
DLM-PromptEngg
This is a modified version of DLM where within the reflection prompt, we include examples for singular queries of how the LLM should reason over the different reward function choices (see Appendix E).
6.3 Results: Composite Prioritization Prompts
We analyze the performance on the $12$ composite prompts which request the prioritization of two subpopulations (see Section D.1 for additional results).
@@ Results
Table 1: Results comparing different methods for aligning with prioritization prompts while minimizing unintended utility shifts. A higher summed utility change implies a better alignment, whereas less unintended shift is better.
Synthetic Domain Real-World Domain
Summed % Utility Change
in Desired Feature(s)
@@
Summed % Utility Change
in Desired Feature(s)
@@
DLM-OP -0.6$\pm$0.58 2.315$\pm$0.05 2.016$\pm$1.39 0.276$\pm$0.03
DLM-EP -1.882$\pm$0.87 2.346$\pm$0.08 -0.254$\pm$1.45 0.276$\pm$0.04
SCLM 4.854$\pm$0.55 0.12$\pm$0.01 6.01$\pm$1.8 0.062$\pm$0.02
6.4 Results: Utility Shifts & Cumulative Utility
@@
As discussed in Section 5.2, we can also use our pipeline to prevent unintended side-effects of aligning reward functions with prioritization clauses, i.e., (i) shifts in the utility feature distribution of features not included in the prompt and (ii) drops in the total generated utility. We focus on (a) here and relegate the results for (b), which paint a very similar picture, to Section D.3. We analyze all $6$ singular and $12$ composite prompts, where we add additional clauses to prevent shifts in the utility distribution of all features not referenced in the prompt. We use SCLM-SIM-Utilitarian as our method. As baselines, we consider DLM only prompted with the prioritization clause(s) (called DLM-OP) and DLM prompted with the prioritization clause(s) and clause(s) asking for the minimizing of utility shifts (called DLM-EP).
@@
To compute the alignment with the prioritization clauses, similar to Section 6.3, we compute the average change in the utility generated by the prioritized subpopulations. To quantify the unintended utility shifts, we compute the average Earth mover’s distance between the utility feature distribution under the candidate and default reward function for each feature not included in one of the prioritization clauses.
@@
Table 1 shows the results. Comparing DLM-OP and DLM-EP, we find that adding the additional objective to the prompt does not result in a better performance for the synthetic and real-world domains. In contrast, SCLM chooses reward functions resulting in significantly higher utility increases for the prioritized subpopulations and significantly fewer unintended utility shifts. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types).
7 Discussion
@@
We present a customizable Social Choice Language Model to handle the multi-objective nature of preference prompts in reward design for RMABs. We showcase how methods from social choice can be used to improve the quality and transparency of decision-making of LLM-based frameworks, as we present an adjudicator component that makes the final decision from options generated by the LLM. SCLM significantly improves the quality of the chosen reward functions. We demonstrate that SCLM can not only handle composite prioritization prompts but arbitrary prompts containing multiple objectives, e.g., balancing the prioritization of subpopulations with the total utility generated by all agents. For future work, SCLM can be applied to other problems from multiagent planning and reinforcement learning. Further, SCLM can easily be extended to handle multiple preference prompts specified by different users.
@@ Appendix A Additional Related Work
Real-world decision-making often involves trade-offs between multiple, conflicting objectives, making multi-objective reinforcement learning (MORL) essential. A common approach in MORL is to scalarize multiple objectives into a single reward function, but this can oversimplify the problem. To address this, Pareto-based methods identify Pareto-optimal solutions that reflect different trade-offs between objectives, providing a more comprehensive decision-making framework (Roijers et al. 2013; Van Moffaert and Nowé 2014). Recent advancements in MORL have introduced algorithms that enhance scalability, handle non-linear objectives, and improve exploration strategies. For example, (Yang, Sun, and Narasimhan 2019) proposed an algorithm that adapts policies based on changing objective importance. Additionally, empirical methods have been developed to evaluate MORL performance across scenarios (Vamplew et al. 2011).
In multi-agent systems, MORL adds complexity as agents must balance their objectives while considering others’. Specialized approaches, such as extending Pareto-based methods to multi-agent contexts, aim to find Pareto-efficient solutions for all agents involved (Bloembergen et al. 2015). Moreover, the concept of fairness becomes important in multi-agent MORL, as solutions must be equitable to all agents while still optimizing overall system performance. Researchers have begun to investigate fairness-aware MORL algorithms that explicitly consider the distribution of rewards among agents, ensuring that no single agent disproportionately benefits at the expense of others (Jiang and Lu 2019; Zimmer et al. 2021).
@@
Designing effective reward functions for RL has long been a challenge due to the lack of robust learning signals. (Ladosz et al. 2022). Previous approaches for reward design in RL include manually specifying reward components and tuning their coefficients (Ng, Harada, and Russell 1999; Margolis and Agrawal 2023), and learning reward models from demonstrations (Ho and Ermon 2016) or preferences (Zhang and Ramponi 2023). While promising, these methods often demand substantial domain expertise, large amounts of data, or computationally expensive processes.
@@
In our adjudicator, we take a social choice perspective to model the problem of selecting a final reward function in a way that balances different objectives against each other. Principles from social choice have informed LLM-based algorithm in previous works, albeit in the very different context of creating a consensus statements for a group of users with diverse views (Bakker et al. 2022; Fish et al. 2023). Moreover, social choice principles are also increasingly used to inform the fine-tuning of language models, including its alignment with human feedback (Bakker et al. 2022; Conitzer et al. 2024; Ge et al. 2024; Dai and Fleisig 2024).
Appendix B Additional Material for Section 4
(a) Income feature.
(b) Education feature.
Figure 4: Utility feature distributions for the default reward function (orange) and reward function returned for prompt “Prioritize agents with low income” (blue). The $x$-axis depicts the values of the feature and the $y$-axis the total utility generated by all agents with this feature value.
Unintended Utility Shifts
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this might change might lead to unintended utility shifts for features not specified in the prompt. In Figure 4, we show an instance of our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “ Prioritize agents with low income ” (orange) and the default reward (blue). While the utility generated by low-income agents increases, the utility generated by highly educated agents decreases, a side-effect a user is likely unaware of and that might conflict with their allocation goals.
Appendix C Implementation Details
C.1 Whittle Index Policy

### v3

@@ 5.2. Adjudicator
Social welfare functions select an alternative based on input preferences of voters. The pros and cons of individual social welfare functions have been extensively researched and debated in social choice Arrow et al. (2010); Rawls (2017). We make use of cardinal social welfare functions Keeney and Kirkwood (1975) which take as input our alignment scores $(s_{i}(R))_{i\in[\ell],R\in\mathcal{R}}$ and output the winning reward function. We consider the three arguably most popular social welfare functions:
Utilitarian:
@@
For each preference clause $p_{i}\in P$, we compute a reward function $R_{i}$ aligned with $p_{i}$ by casting it as a singular prompt to the DLM pipeline (see Section B.2). For each $R\in\mathcal{R}$, we compute as $s_{i}(R)$ the expected reward according to reward function $R_{i}$ produced by policy $\Pi(R)$ (again, we approximate this quantity by running multiple simulations). $s_{i}(R)$ quantifies the quality of the policy induced by the candidate reward function $R$ from the perspective of $p_{i}$ (as captured by $R_{i}$). As the scale of the reward functions can vary significantly among preference clauses, we normalize the scores by the performance of the default policy, i.e., we compute $\frac{s_{i}(R)-s_{i}(R^{*})}{s_{i}(R^{*})}$.
@@
The Simulator Scorer Model assumes access to reward functions capturing individual preference clauses well. If no well-aligned reward functions can be obtained, the performance of SCLM-SIM can deteriorate because it can become very noisy. Another disadvantage of SCLM-SIM is that the scores in SCLM-SIM are all computed via simulation, which can become computationally quite costly. Motivated by this, we propose a quicker and more flexible LLM-based approach, where we prompt an LLM to rate the alignment of a candidate reward function with a preference clause. In particular, for each $R\in\mathcal{R}$ and $p_{i}\in P$, we use a prompt that includes $R$, $p_{i}$, and the utility feature distributions produced by policy $\Pi(R)$. We ask the LLM to rank how well $R$ aligns with the preference clause $p_{i}$ on a scale from 1 to 5 (see Appendix D for prompt texts).
@@
Aligning reward functions to a prioritization prompt may cause (unintended) utility shifts in other features (e.g., due to feature correlations, shifting utility to low-income beneficiaries might shift it away from more educated ones). See Figure 2 for a concrete example. SCLM offers users the option to explicitly prevent these shifts by adding additional clauses (“objectives”) to the prompt: Given a prompt $P$ (e.g., the prompt from Section 4), for each feature not referenced in the prompt, the user can add a new preference clause requesting a minimum shift in the utility distribution of this feature (e.g., for Section 4 they could add “do not change the utility distribution for feature $C$”). To compute the alignment score $s_{i}(R)$ between a reward function $R$ and a clause $p_{i}$=“minimize utility shift for feature $X$”, we compare feature $X$’s utility distribution under the default policy with its utility distribution under the policy $\Pi(R)$. Specifically, we quantify the difference using the Earth mover’s distance (EMD) between the two distributions. Afterward, we apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$, which are input to the social welfare function (along with the alignment scores for the other clauses).
@@
6. Experiments
We describe our testset (Section 6.1), the compared methods (Section 6.2), and our experimental results both for dealing with composite prioritization prompts (Section 7.1) and additionally minimizing unintended side effects (Section 5.2). Following the work of Behari et al. (2024), which constitutes our most important baseline, we use Gemini Pro Anil et al. (2023) as the LLM in our experiments.
6.1. Dataset Description
@@
We create three synthetic datasets, each with three features ($A$, $B$, and $C$). For some agent and feature, we randomly sample the agent’s feature value between 00 and $1$. Arm’s passive transition probabilities ($P(s,a=0,s^{\prime})$ for $s,s^{\prime}\in[0,1]$) are sampled uniformly between 00 and $1$. Our three datasets differ in how active transition probabilities are sampled. For each dataset, we set a three-dimensional weight vector $W\in[0,1]^{3}$ specifying how much each feature impacts the effect of applying an active action (see Section B.3 for details). For each agent, let $f$ denote their feature values; we sample their active transition probabilities for $s,s^{\prime}\in[0,1]$ as $P(s,a=1,s^{\prime})=P(s,a=0,s^{\prime})+\delta,\;\delta\sim\mathcal{N}(\Delta,\sigma),$ where $\Delta=W\cdot f^{T}$ and $\sigma$ is a hyperparameter. Subsequently, we discretize the values of all features into $5$ different equal-sized buckets to be consistent with the real-world domain. For each dataset, we sample $5$ instances according to the above procedure with $N=2100$ arms, a budget of $B=210$, and a time horizon of $T=12$ (to replicate the setup of the real-world domain described below where roughly $3\%-10\%$ of beneficiaries receive a resource every week for up to three months).
@@
ARMMAN (2024) is a non-profit in India that operates large-scale Maternal and Child Care Mobile Health programs for underserved communities. One of their programs disseminates critical health information via weekly automated voice messages. The goal of the NGO is to maximize beneficiaries’ engagement, i.e., the number of messages they listen to. A limited number of beneficiaries are called by health workers every week to boost engagement. The problem of planning which beneficiaries to call has been modeled and solved as an RMAB, where the good/bad state corresponds to a high/low weekly engagement of the beneficiary. We use anonymized data from a quality improvement study conducted in January 2022 Verma et al. (2023). For each beneficiary, we have access to their income, education, and age level, which we use as our three features. Beneficiaries’ historic listenership values are used to estimate their transition probabilities under the passive action Mate et al. (2022). One problem in estimating transition probabilities under the active action is that due to the limited number of service calls made, such transitions are rare. Thus, active transition probability estimates are noisy. To alleviate this issue, we use the features and passive transition probabilities from ARMMAN together with the procedure to sample active transitions described in the synthetic domain. Again, we create datasets for three different weight vectors $W$. Each dataset consists of five sampled RMAB instances with $N=2100$ arms, a budget of $B=210$ and a time horizon of $T=12$.
@@
6.2. Models & Baselines
We analyze six different variants of SCLM differing in the used social welfare function (Utilitarian, Egalitarian, Nash) and scorer model (Simulator or LLM), e.g., we denote as SCLM-SIM-Egalitarian SCLM with the Simulator Scorer Model and the Egalitarian social welfare function. In our generator, we generate $4$ candidate reward functions in each step and run $5$ iterations to generate a total of $20$ candidate reward functions. In addition, we consider several LLM-focused baselines (see Appendix D for detailed descriptions):
LLM-Zeroshot:
@@ 6.2. Models & Baselines
DLM:
This baseline implements the Decision-Language Model by Behari et al. (2024) (see Section B.2).
DLM-PromptEngg:
This is a modified version of DLM where within the reflection prompt, we include examples for singular queries of how the LLM should reason over the different reward function choices (see Appendix D).
7. Results
7.1. Composite Prioritization Prompts
We analyze the performance on the $12$ composite prompts which request the prioritization of two subpopulations (see Section C.1 for additional results).
@@ 7.1. Composite Prioritization Prompts
Figure 4. Results comparing the quality of reward design methods for composite prioritization prompts. Results are averaged across $180=12\cdot 15$ values: $12$ composite prompts on $15$ RMAB instances (from $3$ datasets). Error bars represent std-error.
@@ Results
Effect of Social Welfare Function
Figure 5. Comparison of different methods using the prompt “Prioritize agents with a low value for feature $B$ and agents with a low value for feature $C$”.
While both the Utilitarian and Egalitarian social welfare function leads to good overall results, the choice between them significantly influences which reward function is selected on the instance level. Figure 5 shows one example instance for a prompt with two prioritization clauses. As in Figure 1, each point corresponds to a candidate reward function with the axes quantifying the alignment with the two respective clauses (as described in the beginning of this section). The SCLM-Utilitarian model chooses a reward function from the Pareto frontier that shows a much stronger effect for the second clause (i.e., the utility increase for agents with a low value of feature $C$ is more pronounced). In contrast, the reward function selected by SCLM-Egalitarian is much more balanced in distributing the utility increase between the two groups. DLM-based baselines fail to pick a reward function from the Pareto frontier.
Table 1. Results comparing different reward function selection strategies, aggregated across the three real-world datasets. A higher summed % change in desired feature(s) implies a better alignment with the prioritization clauses, whereas less unintended shift and less % drop in utility are better.
Method
Summed % Change in
Desired Feature(s)
@@
DLM-PrioritizationOnly 6.809$\pm$0.86 0.302$\pm$0.02
DLM-ExtendedPrompt-Fair -0.254$\pm$0.73 0.276$\pm$0.02
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.316$\pm$0.02
SCLM-ExtendedPrompt-Fair 15.364$\pm$0.94 0.099$\pm$0.01
(a) Tradeoff between prioritization prompts alignment and unintended shift minimizationy
Method
Summed % Change in
Desired Feature(s)
@@
% Drop in
Utility
DLM-PrioritizationOnly 6.809$\pm$0.86 0.302$\pm$0.02 6.233$\pm$0.17
DLM-ExtendedPrompt-Util 3.416$\pm$1.04 0.241$\pm$0.02 5.689$\pm$0.25
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.316$\pm$0.02 5.372$\pm$0.19
SCLM-ExtendedPrompt-Util 8.869$\pm$0.73 0.323$\pm$0.02 3.91$\pm$0.21
SCLM-Full 11.496$\pm$0.87 0.142$\pm$0.01 3.85$\pm$0.18
(b) Tradeoff between prioritization prompts alignment and total utility maximization
7.2. Addressing Fairness and Biases
@@
As discussed in Section 5.2, we can also use our pipeline to prevent unintended side-effects of aligning reward functions with prioritization clauses: In this section, we will focus on preventing shifts in utility feature distribution of features not included in the prompt.
We analyze all $6$ singular and $12$ composite prioritization prompts (see Section 6.1), where we add additional clauses to prevent shifts in the utility distribution of all features not referenced in the prompt. We use the simulator scorer with a utilitarian social welfare function and call the resulting model SCLM-ExtendedPrompt-Fair. As baselines, we consider DLM only prompted with the prioritization clause(s) (called DLM-PrioritizationOnly) and DLM prompted with the prioritization clause(s) and clause(s) asking for the minimizing of utility shifts for the other features (called DLM-ExtendedPrompt-Fair). We also consider SCLM-SIM-Utilaterian only prompted with the prioritization clause(s) (called SCLM-PrioritizationOnly). See Appendix D for more details of prompts.
@@
To compute the alignment with prioritization clauses, similar to Section 7.1, we compute average change in utility generated by prioritized subpopulations. To quantify unintended utility shifts, we compute the average Earth mover’s distance between utility feature distribution under the candidate and default reward function for each feature not included in one of the prioritization clauses.
@@
Figure 5(a) shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Fair, we find that adding additional objective to the prompt does not result in a better performance for real-world domains. In contrast, SCLM-ExtendedPrompt-Fair which incorporates unintended shifts in the selection chooses reward functions resulting in significantly higher utility increases for prioritized subpopulations and significantly fewer unintended utility shifts. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types). We see similar results in synthetic domain (see Table 5 in the appendix).
7.3. Preventing Ineffective Resource Use
We additionally consider situations where the user specifies prioritization prompts for multiple attributes but is still interested in maximizing the total generated utility. As discussed at the end of Section 5.2, our framework can handle such settings by adding utility maximization as an additional evaluation dimension. For all 6 singular and 12 composite prioritization prompts in our tests, we add additional clauses to maximize total utility and call the resulting model SCLM-ExtendedPrompt-Util. As a baseline, we consider three models: DLM/SCLM prompted only by the prioritization clause (DLM/SCLM-Prioritization Only) and DLM prompted by both the prioritization clause(s) and a clause requesting the maximization of overall utility (DLM-ExtendedPrompt-Util).
As before, to assess alignment with prioritization clauses, we calculate the average change in utility for the prioritized subgroup. To quantify the performance of a reward function with regards to maximizing utility, we calculate the percent drop in utility as compared to the default reward function.
Figure 5(b) shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Util, we again observe that adding the additional objective to the prompt does not result in a better performance in real-world domains (for similar results on the synthetic domain, see Table 6 in the appendix). In contrast, SCLM-ExtendedPrompt-Util which incorporates utility as a scoring dimension chooses reward functions resulting in a significantly higher utility increases for the prioritized subpopulations and a significantly higher total utility (i.e., there is a lower drop in utility).
Combining Prioritization, Unintended Shifts, and Utility Drops
Finally, we consider a variant of SCLM which both prevents unintended shifts and utility drops, i.e., in addition to the specified prioritization clauses, we add clause(s) on the minimization of shifts and a clause on the minimization of utility drops (which results in having a total of 4 evaluation dimensions; one per feature plus the utility drop one). We use the simulator scorer together with the utilitarian welfare function and call the resulting model SCLM-Full. We see in the last row of Figure 5(b) that this model chooses balanced reward functions with high % change in utility in desired attributes, high overall utility, and low unintended shift, showcasing that SCLM can flexibly allow the end-user to combine multiple evaluation dimensions.
8. Discussion
@@
We present a customizable Social Choice Language Model to handle the multi-objective nature of preference prompts in reward design for RMABs. We showcase how methods from social choice can be used to improve the quality and transparency of decision-making of LLM-based frameworks, as we present an adjudicator component that makes the final decision from options generated by the LLM. SCLM significantly improves the quality of the chosen reward functions. We demonstrate that SCLM can not only handle composite prioritization prompts but arbitrary prompts containing multiple objectives, e.g., balancing the prioritization of subpopulations with the total utility generated by all agents. For future work, SCLM can be applied to other problems from multiagent planning and reinforcement learning. Further, SCLM can easily be extended to handle multiple preference prompts specified by different users. As our model is compatible with any social welfare function, it would also be possible to let the user manually configure the social welfare function. One could in particular allow for adding weights in front of the different alignment scores, enabling the user to express the importance of different clauses (respectively, objectives).
@@ Appendix A Additional Related Work
Real-world decision-making often involves trade-offs between multiple, conflicting objectives, making multi-objective reinforcement learning (MORL) essential. A common approach in MORL is to scalarize multiple objectives into a single reward function, but this can oversimplify the problem. To address this, Pareto-based methods identify Pareto-optimal solutions that reflect different trade-offs between objectives, providing a more comprehensive decision-making framework Roijers et al. (2013); Van Moffaert and Nowé (2014). Recent advancements in MORL have introduced algorithms that enhance scalability, handle non-linear objectives, and improve exploration strategies. For example, Yang et al. (2019) proposed an algorithm that adapts policies based on changing objective importance. Additionally, empirical methods have been developed to evaluate MORL performance across scenarios Vamplew et al. (2011).
In multi-agent systems, MORL adds complexity as agents must balance their objectives while considering others’. Specialized approaches, such as extending Pareto-based methods to multi-agent contexts, aim to find Pareto-efficient solutions for all agents involved Bloembergen et al. (2015). Moreover, the concept of fairness becomes important in multi-agent MORL, as solutions must be equitable to all agents while still optimizing overall system performance. Researchers have begun to investigate fairness-aware MORL algorithms that explicitly consider the distribution of rewards among agents, ensuring that no single agent disproportionately benefits at the expense of others Jiang and Lu (2019); Zimmer et al. (2021).
@@
Designing effective reward functions for RL has long been a challenge due to the lack of robust learning signals. Ladosz et al. (2022). Previous approaches for reward design in RL include manually specifying reward components and tuning their coefficients Ng et al. (1999); Margolis and Agrawal (2023), and learning reward models from demonstrations Ho and Ermon (2016) or preferences Zhang and Ramponi (2023). While promising, these methods often demand substantial domain expertise, large amounts of data, or computationally expensive processes.
@@
In our adjudicator, we take a social choice perspective to model the problem of selecting a final reward function in a way that balances different objectives against each other. Principles from social choice have informed LLM-based algorithm in previous works, albeit in the very different context of creating a consensus statements for a group of users with diverse views Bakker et al. (2022); Fish et al. (2023). Moreover, social choice principles are also increasingly used to inform the fine-tuning of language models, including its alignment with human feedback Bakker et al. (2022); Conitzer et al. (2024); Ge et al. (2024); Dai and Fleisig (2024).
Appendix B Implementation Details
B.1. Whittle Index Policy